# The Responsible Development of Automated Student Feedback with Generative AI

### v2

@@ Abstract
Contribution: This paper identifies four critical ethical considerations for implementing generative AI tools to provide automated feedback to students.
Background: Providing rich feedback to students is essential for supporting student learning. Recent advances in generative AI, particularly with large language models (LLMs), provide the opportunity to deliver repeatable, scalable and instant automatically generated feedback to students, making abundant a previously scarce and expensive learning resource. Such an approach is feasible from a technical perspective due to these recent advances in Artificial Intelligence (AI) and Natural Language Processing (NLP); while the potential upside is a strong motivator, doing so introduces a range of potential ethical issues that must be considered as we apply these technologies.
Intended Outcomes: The goal of this work is to enable the use of AI systems to automate mundane assessment and feedback tasks, without introducing a “tyranny of the majority”, where the needs of minorities in the long tail are overlooked because they are difficult to automate.
Application Design: This paper applies an extant ethical framework used for AI and machine learning to the specific challenge of providing automated feedback to student engineers. The task is considered from both a development and maintenance perspective, considering how automated feedback tools will evolve and be used over time.
Findings: This paper identifies four key ethical considerations for the implementation of automated feedback for students: Participation, Development, Impact on Learning and Evolution over Time.
@@ I Introduction
While assessment is a clear space of development for this type of educational technology, we argue that the real potential of generative language modelling can be found in student feedback. We propose that this type of technology can, relatively straightforwardly, be developed such that a student in any educational program can essentially receive unbounded amounts of feedback. One advantage of this approach over current approaches is that the existing feedback mechanisms using, e.g., multiple choice questions or parametrized questions in engineering courses are limited by several factors. For one, such methods do not easily scale, and tend to tie educational programs into a strict set of practices, where it is difficult to argue for innovation due to the prohibitive costs of developing new feedback scripts. Although many advantages of automation are gained from these practices, e.g., recognize common patterns of student answers and standardize responses to them, rather than having to make bespoke responses to them all, there is a further weakness in that this feedback does not extend to a large portion of students. We argue that current approaches to automated feedback best serve the median student, as developing tailored feedback for the long tail of students—be it highly excelling students who can be helped to excel further, or struggling students who need specific help to succeed in their education—is more expensive, both in terms of human and fiscal resources, and helps fewer students. The flexibility and scalability of generative language modelling provide a unique solution to this tyranny of the long tail(s). This, in turn, allows us to consider real universal coverage using automated feedback, allowing us to meet our obligations to all students, not just those who present the most common submissions.
To move towards this goal, we must consider how we generate this feedback, who engages with this feedback (and when and how), and the impact of providing feedback in a new paradigm. This requires the development of a framework for responsible development, where we follow recommendations from previous work in that we look at the dangers of automation in engineering education, aiming to keep engineering education human-centered [7, 8]. While doing so we must avoid the “Turing Trap” (i.e., realistic human intelligence automation via an excessive focus on automation based on AI technologies) [9]. The development of this framework denotes the core contribution of this article, in which we:
@@ III The Ethical Issues Involved in the Provision of Automated Feedback
These dimensions guide the application of generative AI for automated feedback, ensuring all students benefit. By these means, we propose the following four ethical questions with elaboration surrounding the provision of automated feedback and should be taken into consideration for developing these systems in the future. They fall into four key categories:
@@ III-B Development
III-C Impact on Learning
@@ Acknowledgments
This research is supported by the research grant (VIL57392) from VILLUM FONDEN.

### v3

@@ Abstract
Providing rich, constructive feedback to students is essential for supporting and enhancing their learning. Recent advancements in Generative Artificial Intelligence (AI), particularly with large language models (LLMs), present new opportunities to deliver scalable, repeatable, and instant feedback, effectively making abundant a resource that has historically been scarce and costly. From a technical perspective, this approach is now feasible due to breakthroughs in AI and Natural Language Processing (NLP). While the potential educational benefits are compelling, implementing these technologies also introduces a host of ethical considerations that must be thoughtfully addressed.
One of the core advantages of AI systems is their ability to automate routine and mundane tasks, potentially freeing up human educators for more nuanced work. However, the ease of automation risks a “tyranny of the majority”, where the diverse needs of minority or unique learners are overlooked, as they may be harder to systematize and less straightforward to accommodate. Ensuring inclusivity and equity in AI-generated feedback, therefore, becomes a critical aspect of responsible AI implementation in education.
The process of developing machine learning models that produce valuable, personalized, and authentic feedback also requires significant input from human domain experts. Decisions around whose expertise is incorporated, how it is captured, and when it is applied have profound implications for the relevance and quality of the resulting feedback. Additionally, the maintenance and continuous refinement of these models are necessary to adapt feedback to evolving contextual, theoretical, and student-related factors. Without ongoing adaptation, feedback risks becoming obsolete or mismatched with the current needs of diverse student populations. Addressing these challenges is essential not only for ethical integrity but also for building the operational trust needed to integrate AI-driven systems as valuable tools in contemporary education. Thoughtful planning and deliberate choices are needed to ensure that these solutions truly benefit all students, allowing AI to support an inclusive and dynamic learning environment.
@@ I Introduction
While assessment is a clear space of development for this type of educational technology, we argue that the real potential of generative language modelling can be found in student feedback. We propose that this type of technology can, relatively straightforwardly, be developed such that a student in any educational program can essentially receive unbounded amounts of feedback. One advantage of this approach over current approaches is that the existing feedback mechanisms using, e.g., multiple choice questions or parametrized questions in engineering courses are limited by several factors. For one, such methods do not easily scale, and tend to tie educational programs into a strict set of practices, where it is difficult to argue for innovation due to the prohibitive costs of developing new feedback scripts. Although many advantages of automation are gained from these practices, e.g., recognize common patterns of student answers and standardize responses to them, rather than having to make bespoke responses to them all, there is a further weakness in that this feedback does not extend to a large portion of students. We argue that current approaches to automated feedback best serve the median student, as developing tailored feedback for the long tail of students—be it highly excelling students who can be helped to excel further, or struggling students who need specific help to succeed in their education—is more expensive, both in terms of human and fiscal resources, and helps fewer students. The flexibility and scalability of generative language modelling provide a unique solution to this tyranny of the long tail(s). This, in turn, allows us to consider real universal coverage using automated feedback, allowing us to meet our obligations to all students, not just those who present the most common submissions. Although numerous successful case studies (e.g., [7]) highlight ChatGPT’s potential for fostering, e.g., creative problem-solving, the ethical implications of integrating such systems into classroom settings have yet to be thoroughly examined.
To move towards this goal, we must consider how we generate this feedback, who engages with this feedback (and when and how), and the impact of providing feedback in a new paradigm. This requires the development of a framework for responsible development, where we follow recommendations from previous work in that we look at the dangers of automation in engineering education, aiming to keep engineering education human-centered [8, 9]. While doing so we must avoid the “Turing Trap” (i.e., realistic human intelligence automation via an excessive focus on automation based on AI technologies) [10]. In Fig. 2, we visualize our contributions. Overall, the development of this framework denotes the core contribution of this article, in which we:
Figure 2: We explore critical ethical considerations for using generative AI to deliver automated student feedback in education. By applying an ethical framework for AI and machine learning [11, 12], this study addresses both development and maintenance phases, emphasizing how these tools will evolve over time. Key ethical issues highlighted include inclusive participation, responsible development, impact on student learning, and adaptability over time.
@@ III The Ethical Issues Involved in the Provision of Automated Feedback
These dimensions guide the application of generative AI for automated feedback, ensuring all students benefit. By these means, we propose the following four ethical questions with elaboration surrounding the provision of automated feedback and should be taken into consideration for developing these systems in the future. They fall into six key categories:
@@
  5. 5.
    Scaling Up: Automated feedback systems offer the opportunity of working at scale—how do we account for the broader scope?
  6. 6.
    AI Literacy: To what extent will we rely upon the skills of the users of these systems?
@@ III-B Development
III-C Impact of Automated Feedback on Learning
@@ III-D Evolution over Time
III-E Scaling Up
Maintaining ethical oversight in AI systems is a challenging problem, especially as they scale-up [49]. Scaling up can lead to lack of insights into who is using the system and for what purpose. Scaling is also problematic because a small problem within the system can have a dramatically high impact as a larger number of users are affected [50]. To ethically scale-up assessment systems, it is important to first and foremost test any application before implementing it. Rigorous testing and evaluation are the cornerstone of quality software systems, and AI-based assessment applications are no exception. In this case, though, given the challenges with bias and fairness, testing and evaluation should involve a diverse set of users. Providing good documentation and ensuring transparency and explainability can further support this process by building trust among users and allowing them to understand how decisions are made and have confidence in reporting any misgivings and problems as soon as they occur. This can also help build guidance and guardrails around the acceptable use of an application [51]. It is important then to conduct regular audits of the systems in addition to a process to continuously monitor the application. If needed, third-party audits can also be included in the process. Finally, it is important to remember throughout this process that the application, as well as its ethical oversight, should complement rather than replace human educators. Given the goal of education to support learning by fostering critical thinking and advancing the personal development of learners, it is important that educators are part of the process throughout.
III-F AI Literacy
Finally, one of the critical assumptions made here is that not only those who are building AI-based assessment applications, but even those who are using it, will be AI literate. Given the fast-changing AI environment, this is an assumption as best, and to ensure the best use of AI systems, it is important that educators are AI literate. As the use of generative AI applications has indicated, educators are often far behind students in the adoption of AI and often lack a fundamental understanding of how AI works [52]. Without this knowledge, interventions based on AI can never be fully adopted within educational settings. Therefore, training for educators is essential for proper use of AI systems, especially in an ethical and responsible manner. Faculty need to understand both the technical aspects of how a system work, to some extent, and how the overall application fits within their ecosystem [53]. And even if they do not develop and understanding of how to create such a application, knowing how it works is critical in assessing its output and usefulness for the task at hand [54].
@@ Acknowledgments
EDL, MZ, and JB are supported by the research grant (VIL57392) from VILLUM FONDEN. AJ is partly supported by US NSF Awards 2319137, 1954556, and USDA/NIFA Award 2021-67021-35329. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies.